### v2

# TiDES: The 4MOST Time Domain Extragalactic Survey

Abstract
The Time Domain Extragalactic Survey (TiDES) conducted on the 4-metre Multi-Object Spectroscopic Telescope (4MOST) will perform spectroscopic follow-up of extragalactic transients discovered in the era of the NSF-DOE Vera C. Rubin Observatory. TiDES will conduct a 5-year survey, covering , and use around 250 000 fibre hours to address three main science goals: (i) spectroscopic observations of 30 000 live transients, (ii) comprehensive follow-up of 200 000 host galaxies to obtain redshift measurements, and (iii) repeat spectroscopic observations of Active Galactic Nuclei to enable reverberation mapping studies. The live spectra from TiDES will be used to reveal the diversity and astrophysics of both normal and exotic supernovae across the luminosity-timescale plane. The extensive host-galaxy redshift campaign will allow exploitation of the larger sample of supernovae and improve photometric classification, providing the largest-ever sample of type Ia supernovae, capable of a sub-2 per cent measurement of the equation-of-state of dark energy. Finally, the TiDES reverberation mapping experiment of 700–1 000 AGN will complement the SN Ia sample and extend the Hubble diagram to .
show]c.frohmaier@soton.ac.uk
1 Introduction
The NSF-DOE Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ž. Ivezić et al., 2019) heralds a new era of transient discovery: LSST will produce millions of transient alerts and photometric data for hundreds-of-thousands of supernovae (SNe) and other variable sources. While this discovery rate is exciting, the ability to obtain spectroscopic follow-up of these transients must similarly scale to enable much of the scientific exploitation. Follow-up spectra are critical for extracting the full astrophysical potential of the objects discovered, for example their classifications, rates, chemistry, redshifts, luminosities – and ultimately their physical nature. The capacity to assemble these spectroscopic samples presents a new challenge for the transient community, and to meet this, a next generation facility is needed.
In this paper we present the Time Domain Extragalactic Survey (TiDES), which is designed to meet the spectroscopic requirements of the LSST-era for optical transients, their host galaxies, and active galactic nuclei (AGN). TiDES will use the 4-metre Multi-Object Spectroscopic Telescope (4MOST; R. S. de Jong et al., 2019), a fibre-fed spectroscopic survey facility on the European Southern Observatory (ESO) Visible and Infrared Survey Telescope for Astronomy (VISTA). This facility packs science fibres distributed over the 4.2 deg2 field-of-view, with each spine configurable by the AESOP fibre positioning system onto a target of interest. Throughout this work we refer to the concept of a fibre-hour. A fibre-hour is one hour of observing time, including overheads, for one fibre. In total, 4MOST provides up to 2,436 fibre-hours per one hour of observation. TiDES will exploit the fact that wherever 4MOST points over deg2 of the extragalactic sky, there will be known time-variable sources recently discovered by Rubin or other surveys. These will include recently discovered transients, and older, faded events for which host galaxy spectroscopy will be obtained for measuring redshifts. The strategy takes inspiration from the successful ‘OzDES’ survey (F. Yuan et al., 2015; C. Lidman et al., 2020), conducted using the 2dF fibre positioner and the AAOmega spectrograph on the Anglo-Australian Telescope, which secured live SN spectra and 4,500 redshifts (M. Vincenzi et al., 2019) for SN host galaxies over 27 deg2 of imaging from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016).
TiDES will use around 250 000 fibre-hours to tackle three key science goals. The first is the use of type Ia supernovae (SNe Ia) as a cosmological probe. SNe Ia are an exceptionally mature and well-understood cosmological probe via their use as standardisable candles (A. G. Riess et al., 1998; S. Perlmutter et al., 1999; M. Betoule et al., 2014; DES Collaboration et al., 2024). They remain a uniquely powerful distance-indicator in the Universe, and directly constrain the properties of dark energy when combined with Planck Cosmic Microwave Background (CMB) (Planck Collaboration et al., 2020) and/or Baryon Acoustic Oscillation (BAO; e.g., A. G. Adame et al., 2025) measurements. When combined with these probes (CMB, BAO), current SN samples measure the average dark energy equation-of-state parameter to per cent (D. M. Scolnic et al., 2018; D. Brout et al., 2022; DES Collaboration et al., 2024). This includes a systematic error budget that is comparable to the statistical uncertainty, and show to be broadly consistent with a cosmological constant (), with the intriguing possibility of a time-varying that motivates the need for larger samples of SNe Ia to higher-. Using SNe Ia detected by LSST, TiDES will enable the creation of a Hubble Diagram of cosmologically-useful SNe Ia and make a per cent measurement of , when including a CMB prior, and considering only statistical uncertainties. These constraints are a factor 10 improvement compared to DES-SN5YR. Massive spectroscopic follow-up of SNe Ia will also allow their use in different contexts, for example their brightnesses are lensed by structure, tracing matter distributions along their line of sight (S. Refsdal, 1964; D. E. Holz, 1998; R. B. Metcalf, 1999; J. Jönsson et al., 2010; M. Smith et al., 2014; E. Macaulay et al., 2020; P. Shah et al., 2024a, b), and, with LSST, SNe Ia will be used to trace peculiar velocities on large scales and constrain the growth rate of structure (C. Howlett et al., 2017).
The extragalactic time-domain universe is a far more diverse environment than imagined only a decade ago, and exploring the explosive timescale-luminosity phase-space forms the second goal of TiDES. Objects such as: superluminous supernovae (SLSNe; e.g., R. M. Quimby et al., 2011; C. Inserra et al., 2013; C. R. Angus et al., 2019; T. Kangas et al., 2022; Z. H. Chen et al., 2023, for a review see A. Gal-Yam 2019), calcium-strong (-rich) transients (interchangeably referred to as CaST or CaRT in the literature e.g., M. M. Kasliwal et al., 2012; K. De et al., 2020), exotic thermonuclear explosions (see review by S. Taubenberger, 2017), fast, blue optical transients (FBOTs; e.g., M. R. Drout et al., 2014; M. Pursiainen et al., 2018; S. J. Prentice et al., 2018; R. Margutti et al., 2019; A. Y. Q. Ho et al., 2023), tidal disruption events (TDEs; e.g., M. J. Rees, 1988; S. Gezari et al., 2012; T. W. S. Holoien et al., 2014; S. van Velzen et al., 2021; S. Gezari, 2021), other ambiguous nuclear transients (E. Kankare et al., 2017; B. Trakhtenbrot et al., 2019; P. Wiseman et al., 2023), and kilonovae (N. R. Tanvir et al., 2013; I. Arcavi et al., 2017; D. A. Coulter et al., 2017; V. M. Lipunov et al., 2017; S. J. Smartt et al., 2017; M. Soares-Santos et al., 2017; N. R. Tanvir et al., 2017; S. Valenti et al., 2017; A. J. Levan et al., 2024, for a review see R. Margutti & R. Chornock 2021), have each demonstrated how little is known about explosive transient populations. TiDES will enlarge the population of spectroscopically-confirmed transients by an order of magnitude, and potentially uncover entirely new forms of explosions. This will include 9 000 core collapse SNe, 3 000 SLSNe for probing the universe, and 500 faint-and-fast SN Ia-like events for studying the full range of thermonuclear white dwarf explosions.
The third goal of TiDES is cosmology and galaxy evolution with AGN. These are among the most energetic sources in the Universe, showing variability in all wavebands when mass is accreted onto supermassive black holes in the centres of galaxies. The variability of the optical continuum light from the accreting matter, together with its delayed response mirrored in the broad emission lines of the surrounding material can be used i) to dynamically measure black hole masses (e.g., Y. Shen et al., 2016) and identify sub-parsec-scale supermassive black hole binaries among the AGN population, and ii) as a standardisable candle to high redshifts (R. D. Blandford & C. F. McKee, 1982; B. M. Peterson, 1993; M. Haas et al., 2011; D. Watson et al., 2011; S. Panda & P. Marziani, 2023; B. Czerny et al., 2023a). TiDES will establish an SNe-independent Hubble diagram of AGN and will deliver one of the, if not the, largest catalogue of dynamically measured supermassive black hole masses over redshifts . The expected accuracy of both the Hubble diagram as well as the black hole masses will depend critically on the achieved season lengths and total baseline of the survey (U. Malik et al., 2022) as well as characterisation of the objects in terms of luminosity and accretion rate (e.g. P. Du et al., 2016; G. Fonseca Alvarez et al., 2020; R. Prince et al., 2023; M. Zajaček et al., 2024b).
This paper introduces the TiDES survey and demonstrates its potential using simulations of the LSST and 4MOST survey strategy. In Section 2 we describe the 4MOST spectrograph and the TiDES strategy. Section 3 details our TiDES simulations, and Section 4 presents sample statistics. Our expectation for the SN Ia cosmological analysis is detailed in Section 5. We then summarise in Section 6. Throughout, we assume a baseline cosmology with a matter density (e.g. Planck Collaboration et al., 2020) in a flat universe and a Hubble constant of km s-1 Mpc-1, and use the AB photometric system (J. B. Oke & J. E. Gunn, 1983).
2 4MOST and the TiDES programme
4MOST is a new high-multiplex, wide-field spectroscopic facility under development for VISTA, located on the ‘NTT peak’ at the Cerro Paranal in northern Chile. A full overview can be found in R. S. de Jong et al. (2019), with the scientific operations model described in C. J. Walcher et al. (2019), and the general survey strategy overview in G. Guiglion et al. (2019). 4MOST has a large field-of-view of 4.2 deg2 with a high multiplex capability: 1624 fibres feeding two low-resolution spectrographs (LRS), and 812 fibres feeding a high-resolution spectrograph (HRS). The LRS offers over 3700–9500Å, and the HRS offers in three 400–600Å wide passbands. Within a 1-hour observation, 4MOST has the sensitivity to obtain redshifts of mag galaxies or AGN using the LRS. TiDES will use the LRS fibres.
The 4MOST project as a whole consists of 25 consortium and community surveys each with their specific science goals. Each survey is characterised by a target selection algorithm, but all surveys share the focal plane and are observed in parallel, with individual targets for any given observation selected by a fibre-to-target assignment algorithm (E. Tempel et al., 2020a). It is expected that TiDES will average around 30–35 LRS fibres in each extragalactic 4MOST pointing – around two per cent of the LRS fibres available.
2.1 The TiDES science programme
The TiDES science programme comprises three interlocking surveys each arising from 4MOST’s ability to conduct extensive spectroscopic follow-up of extragalactic transients located with LSST, whilst exploiting additional diagnostic data from other facilities (e.g., Euclid). In this section we describe each of the three surveys.
2.1.1 TiDES-Live
The first survey focuses on the spectroscopic study of live transients (‘TiDES-Live’). This survey aims to observe more than 30 000 live transients discovered by Rubin and other photometric surveys as soon as feasible after discovery. There are two main goals:
1.
A systematic search and classification effort for the full range of extragalactic transients across the luminosity–timescale plane, leading to improved rates and physical understanding of the various events, and
2.
The rapid follow-up of live SN Ia candidates, which will lead to a new Hubble diagram for cosmological studies, while also serving as a critical spectral training set for the larger LSST SN Ia sample studied (and classified) photometrically.
Work over the last decade has charted the luminosity–timescale plane for transients, and uncovered rare and new classes of transients. TiDES-Live will conduct a systematic campaign of follow-up that will be both unbiased and an order of magnitude larger in number than earlier efforts of this nature (e.g., C. Lidman et al., 2020), thereby providing improved rates for both new and well-understood phenomena with full photometric light curves. Early time spectra (3–4 days post-detection) for all classes of SNe will provide new insights into the progenitor environment and the outer layers of the ejecta, and, in 4MOST pointings with repeat visits, multiple spectra will trace SN evolution into the nebular phase. TiDES will also collect the largest spectroscopic sample of rare transients (e.g., calcium-strong fast transients and TDEs) to comprehensively study their important contributions to galactic chemical evolution.
For SN Ia cosmology, TiDES-Live will create an impressive Hubble diagram of more than 10 000 cosmologically-useful events over the redshift range for which there will be associated Rubin deep imaging. A further goal is to explore the dispersion in the Hubble diagram as a function of astrophysical properties (e.g., morphology, SN colour, SN spectral features), with the aim of improving the cosmological precision using SN Ia subsets. Moreover, as even the most advanced machine-learning SN photometric classification techniques are dependent on large, homogeneous and representative training samples (M. Lochner et al., 2016; A. Möller & T. de Boissière, 2020; H. Qu & M. Sako, 2022), the TiDES sample will become an unsurpassed training sample for SN photometric classifiers (J. E. Carrick et al., 2021) unlocking the potential of the entire LSST SN dataset.
2.1.2 TiDES-Hosts
The second survey is spectroscopy of SN host galaxies (‘TiDES-Hosts’). TiDES will secure the spectroscopic redshifts for the host galaxies of SNe of all types for which live SN spectroscopy was not possible, or for both SN and host in regions with dense 4MOST follow up. These redshifts, coupled with photometric classification of the SN events themselves, will form a key component of the LSST SN Ia cosmology programme (The LSST Dark Energy Science Collaboration et al., 2018). 4MOST should secure galaxy redshifts over the range in two hours, and to higher redshifts in fields where multiple repeat exposures can be stacked.
In addition to providing galaxy redshifts essential for the LSST SN Ia Hubble Diagram, TiDES-Hosts will provide further diagnostics of the dispersion in SN Ia magnitudes, such as host galaxy metallicity and star-formation rate (following techniques presented in e.g. J. Dumayne et al., 2023; L. Galbany et al., 2022). Redshifts also enable the intrinsic properties of individual transients, like luminosity and evolution timescale, to be measured. TiDES-Hosts will provide the largest and most precise measure of the cosmological parameters over the key redshift range where dark energy becomes a dominant component.
In this work we consider host galaxies from TiDES in isolation, but it is important to highlight the contribution of other surveys in 4MOST. There are several participating surveys such as the Wide-Area VISTA Extragalactic Survey (WAVES; S. P. Driver et al., 2019) and the 4MOST Hemisphere Survey of the Nearby Universe (4HS; E. N. Taylor et al., 2023) that will observe millions of galaxies over thousands-of-square-degrees. Naturally, some of these galaxies will host supernovae, this will benefit TiDES in two ways: i) TiDES will not need to spend fibre-hours re-observing these galaxies, ii) If 4MOST has completed its planned visits of any given field, redshifts will be available from galaxies hosting future supernovae. These additional observations will supplement the total number of available redshifts for any SN analysis.
2.1.3 TiDES-RM
The final survey concerns repeat spectroscopic observations of AGN for reverberation mapping (‘TiDES-RM’). The survey will focus on long-term variability monitoring of AGN in the deep drilling fields (DDFs) of LSST (A. B. Kovačević et al., 2022; B. Czerny et al., 2023b). Specifically, we will exploit AGN broad line lags to measure dynamical masses of supermassive black holes and construct a Hubble diagram out to redshift , i.e. into cosmic noon. In addition, the continuous high-quality spectroscopic monitoring will enable us to search for any unusual kinematics of the broad lines that could indicate binary supermassive black holes (for a review see A. De Rosa et al., 2019). 4MOST will routinely visit the DDFs every 10-14 days to observe the high density of requested targets and satisfy the deep observing requirements from other surveys. TiDES-RM will use each of the single epoch visits to sample the expected lags of our AGN at a similar rate to past reverberation mapping campaigns (e.g. A. J. Barth et al., 2011). The number of observations required for each object is determined by its lag length, with shorter lags being recovered first. Generally, for a successful recovery of the broad line lags, a single observing season should be longer than a lag length, and an entire observing campaign should be longer than 3 times the lag length.
As standardisable candles, AGN offer a complementary probe to SNe Ia with a reach to higher redshifts and with different astrophysical systematics. Recent work suggests that current probes at low and high redshifts may provide inconsistent results, beyond the well-known tension, pointing either towards unknown systematics in the current methods or new physics (e.g. E. Di Valentino et al., 2019; S. Cao et al., 2022; N. Khadka et al., 2023; S. Cao et al., 2024; M. Zajaček et al., 2024a). TiDES will move forward with a unique test for remaining systematic uncertainties by using two independent standardisable candles, thereby increasing the reliability of the final results. By targeting around 930 AGN over we can improve the precision of constraints based on reverberation mapped AGN by at least a factor of two (A. L. King et al., 2015).
Of equal importance in TiDES-RM will be dynamical mass measures of large numbers of supermassive black holes (SMBH) to , i.e. into cosmic noon, the most active phase of galaxy and black hole growth. Supermassive black hole masses represent a key parameter in governing the assembly history of galaxies through associated feedback processes. Most current SMBH mass measurements outside the local Universe rely on indirect methods, most notably the single-epoch method that takes the width of a broad line and infers the mass via a set of calibrations against local scaling relations. This method is prone to biases and uncertainties as a single emission line width will be unable to capture the dynamical complexities of the underlying accretion and outflow processes and due to the intrinsic difference in AGN demographics at low and high redshift (e.g. F. Shankar et al., 2016; J.-H. Woo et al., 2024). TiDES-RM will measure dynamical SMBH masses at redshifts up to via reverberation-mapping, which is the technique underlying the calibration of local scaling relations (e.g. Y. Shen et al., 2024). In addition, the spatio-kinematic information from emission line lags can be used to trace outliers, with the potential of exposing binary supermassive black holes (e.g. Y.-Y. Songsheng et al., 2019).
3 Simulating LSST+TiDES
We next describe our method for simulating a LSST-like SN survey, and combining with a simulated 4MOST observing strategy. Our simulations focus on LSST for the transient discovery as we anticipate this will provide the majority of the TiDES targets. However, TiDES will be agnostic to the source of the photometric transient feed and will include transients from any source providing a public alert stream e.g. the Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019), the Young Supernova Experiment (YSE; D. O. Jones et al., 2021), BlackGEM (P. J. Groot et al., 2022), The Gravitational-wave Optical Transient Observer (GOTO; D. Steeghs et al., 2022), and The La Silla Schmidt Southern Survey (LS4; A. A. Miller et al., 2025). Thus, this section is intended to produce broad projections for the number and type of transients and hosts that 4MOST may observe, in particular noting that neither LSST nor 4MOST’s strategy are yet finalised. An end-to-end flowchart of our entire simulation framework is shown in Fig. 1.
3.1 Simulating transients
| Transient Class | Description | Source |
| SN Ia | Normal type Ia SNe following SALT2 | R. Hounsell et al. (2018) |
| SN Iax | Type Iax peculiar SNe | R. Kessler et al. (2019a) |
| SN Ia-91bg | Peculiar SNe SN1991bg-like | R. Kessler et al. (2019a) |
| SN II, IIn | Hydrogen-rich core-collapse SNe | M. Vincenzi et al. (2019) |
| SN IIb, Ib, Ic, Ic-BL | Hydrogen-poor stripped envelope SNe | M. Vincenzi et al. (2019) |
| CaRT/CaST | Calcium-rich(-strong) faint and fast SNe | R. Kessler et al. (2019a) |
| SLSN-I | Hydrogen-poor super luminous SNe | M. Nicholl et al. (2017); R. Kessler et al. (2019a) |
| TDE | Tidal disruption events | R. Kessler et al. (2019a); B. Mockler et al. (2019) |
The LSST Science Requirements Document (Ž.. Ivezić & the LSST Science Collaboration, 2013) details the survey specification but with enough room for optimisation of the strategy to balance the different science goals (F. B. Bianco et al., 2022). LSST will have a ‘Wide, Fast, Deep’ (WFD) component coupled with deeper, more frequent observations in ‘Deep Drilling Fields’ (DDFs). Significant work has been performed on the effects of different survey choices for transients (e.g., M. Lochner et al., 2022; C. S. Alves et al., 2023; P. Gris et al., 2023), and we do not attempt to determine an optimal LSST strategy for TiDES in this work. Instead, we use established LSST simulations to create our mock transient datasets, adopting the baseline v3.4111https://survey-strategy.lsst.io/index.html simulation, the latest baseline release at the time of writing. The comparisons with other LSST strategies are presented in the TiDES Cadence Note222 https://docushare.lsst.org/docushare/dsweb/Get/Document-37640/Frohmeier_TiDES.pdf. Note that due to the high cadence of repeat visits in the DDFs, all of our TiDES-RM targets are located there.
The output from the LSST Operations Simulator (OpSim)333https://www.lsst.org/scientists/simulations/opsim is the input to our simulations. Based on a scheduling pattern, OpSim creates a catalogue-level output of the LSST strategy using detailed observing quality metrics based on weather models, maintenance requirements, and telescope hardware constraints. Using these observing characteristics, we then simulate transients and their host galaxies using the SuperNova ANAlysis software snana (R. Kessler et al., 2009) simulation code. snana combines survey-specific metadata (e.g., from OpSim) with the known properties of a SN population to generate a set of light curves as observed by any survey strategy. The OpSim strategies are translated into the appropriate snana format using the OpSimSummary package (R. Biswas et al., 2020). Note that all of our simulations are ‘catalogue level’, i.e., we do not place transients and hosts into simulated images. We assume a saturation limit of 15 mag for the LSST CCD and exclude all points on a simulated light curve exceeding this brightness.
snana generates light curves by integrating a time-evolving spectral energy distribution (‘template’) of a given SN through the survey’s filter-specific response function. Table LABEL:tab:snClasses shows each spectral template we use in our simulations. Our method of generating light curves from spectral templates using snana follows the Photometric LSST Astronomical Time Series Classification Challenge (PLAsTiCC) (R. Kessler et al., 2019a), but we update the core collapse SN templates to those of M. Vincenzi et al. (2019, hereafter V19).
On a technical note, snana precomputes a set of transient sky coordinates with a resolution of several arcminutes. This leads to some simulated events sharing spatial coordinates, which if left unaddressed would lead to artificial fibre collisions in the 4MOST simulator. Increasing this resolution in snana becomes computationally expensive, so we perform a post-processing stage to apply a small coordinate scatter on each transient and host such that their distribution is uniform within an snana resolution element.
3.1.1 Simulation of SNe Ia
We simulate SNe Ia using the SALT2 SED model (J. Guy et al., 2007, 2010) trained following M. Betoule et al. (2014). The SALT2 SED model is described by the SN redshift , the epoch of SN peak brightness , a (dimensionless) stretch-like parameter , and a colour parameter . The SN Ia apparent magnitude in the rest-frame -band is then simulated as
where the parameters and describe the stretch–luminosity and colour–luminosity relations and are set in our simulations to , and . The absolute magnitude for a , SN Ia () is set to -19.365 and is the luminosity distance for redshift with our default cosmological parameters. The term is a randomised brightness offset that encapsulates the intrinsic scatter observed in SN Ia standardised brightness (see J. Guy et al., 2010, for details on the modelling of intrinsic scatter). The absolute number of SN Ia events simulated and their redshift distribution is calculated using the power-law evolution of the volumetric rates presented by C. Frohmaier et al. (2019) (their Equation 6). The intrinsic distributions of SN stretch and colour are taken from D. Scolnic & R. Kessler (2016).
3.1.2 Simulation of other types of transients
All simulated transient types are listed in Table LABEL:tab:snClasses. For core-collapse SNe, we use V19 templates and include both hydrogen-rich core-collapse SN events (type II and IIn SNe) and stripped envelope SNe (Type Ib/IIb/Ic/Ic-BL SNe). For volumetric rates of core-collapse events, we normalise at low redshift to the measurement of C. Frohmaier et al. (2021) and extrapolate to higher redshifts assuming the rate follows the star formation history presented in P. Madau & M. Dickinson (2014). To ensure the correct mix of different core collapse sub-types we use the relative rates presented by I. Shivvers et al. (2017). The intrinsic brightness of simulated events are matched to the luminosity functions presented in W. Li et al. (2011) and revised by V19. Recently, updates have been made to models using different rate assumptions for Type IIn SNe (C. L. Ransome & V. A. Villar, 2025) and templates for stripped-envelope (S. Khakpash et al., 2024), however, these updates do not affect the overall picture for TiDES compared to the models we have used.
We include two types of peculiar SNe Ia: SN1991bg-like events (A. V. Filippenko et al., 1992) and SNe Iax (R. J. Foley et al., 2013). SN1991T-like events are modelled within the normal-SN Ia events in Section 3.1.1. The SED templates and volumetric rates used to simulate these classes of peculiar SNe Ia are those presented by R. Kessler et al. (2019a), applying the revisions in M. Vincenzi et al. (2021, hereafter V21).
Tidal Disruption Events (TDEs), Calcium-Strong Transients (CaST) and hydrogen-poor Superluminous SNe (SLSNe-I) are implemented following the same assumptions and modelling used in the PLASTICC challenge (R. Kessler et al., 2019a).
3.1.3 Simulation of AGN for TiDES-RM
The AGN that we will observe as part of the reverberation mapping survey are already known and are located only in the LSST DDFs. The selection is based on two catalogues, the Million Quasar Catalogue Milliquas (E. W. Flesch, 2023) and the southern photometric quasar catalogue based on Dark Energy Survey DR2 as presented by Q. Yang & Y. Shen (2023). First, both catalogues were cut down to the hexagonal footprints of the 4MOST field-of-view centered on the defined 4MOST deep field positions within the Elais South, COSMOS, ECDF-S, and XMM-LSS deep fields (Table 2). For Milliquas, only sources with a Q or A type flag have been included, which selects against obscured AGN that are unfavourable to reverberation mapping. In a second step, sources from both catalogues were cross-matched. Those present in both catalogues are deemed the “gold sample” while objects only present in Milliquas are part of the “silver sample.” The total number of identified AGN varies with deep field, given how intensely the respective field has been monitored and in which wavebands. Given survey operational constraints, the maximum number of TiDES-RM sources that can be observed per deep field is of the order of 250. To preserve some comparability across the fields, we impose brightness cuts of for the COSMOS and XMM-LSS fields and for the Elais South and ECDF-S fields.
Finally, we want to exclude sources where the survey length will not be able to recover a lag. This requires estimating observed lag length of the cut down catalogues. We use the lag-luminosity relation for the H line presented in M. C. Bentz et al. (2009) and scale to Mg II lags with a factor of 0.9. C IV lags are estimated from the lag-luminosity relation in A. Penton et al. (2022). Approximate luminosities are calculated from band magnitudes. The resulting estimated rest-frame lags have been converted to observed lags by accounting for a cosmological time dilation factor of . We assume that successful lag recovery requires lag lengths to be at least 3 times shorter than the survey length. As such, we cut the target catalogue for AGN where the shortest lag observed in the 4MOST wavelength range for the given redshift is 600 days. With the previously imposed brightness limit, this primarily affects the brightest sources in the redshift range and particularly those at the higher redshift end, where Mg II is the only one of the 3 main target lines present. For all four DDFs, this affects 1% of the pre-cut samples.
| Field | R.A. | Dec. | AGN | H | Mg II | C IV |
| ELAIS-S1 | 9.500 | -43.950 | 191 | 23 | 128 | 80 |
| XMM-LSS | 35.875 | -5.025 | 252 | 47 | 141 | 119 |
| ECDF-S | 53.125 | -28.100 | 230 | 48 | 138 | 92 |
| COSMOS | 150.125 | 2.200 | 256 | 47 | 150 | 94 |
Notes — Redshift ranges covered by individual lines are: H: ; Mg II: ; C IV: . Spectra of some sources will include two broad lines that can be used to cross-calibrate lags: ELAIS-S1: H/Mg II: 10, Mg II/C IV: 30; XMM-LSS: H/Mg II: 17, Mg II/C IV: 38; ECDF-S: H/Mg II: 20, Mg II/C IV: 28; COSMOS: H/Mg II: 14, Mg II/C IV: 21.
The final sample numbers across the four DDFs are listed in Table 2. We highlight number of objects per emission line as well as number of sources in those redshift regions where two lines overlap. These are particularly interesting as they will allow for cross calibration of lag-luminosity relations as well as black hole mass estimates. In total, 61 AGN are expected to have both H and Mg II lags recovered and 117 to overlap in Mg II and C IV.
3.2 Simulating transient host galaxies
SNe Ia, peculiar SNe Ia and core-collapse SNe are associated to host galaxies following the approach and assumptions presented in V21. Host galaxies are selected from a ‘HOSTLIB’ library of 380 000 galaxies from the DES Science Verification galaxy catalogues (A. Drlica-Wagner et al., 2018). Each galaxy in the catalogue has optical photometry from DES () and – when available – near infrared photometry from VISTA. To derive properties of the galaxies, template SEDs were generated using the PÉGASE.2 spectral synthesis code and fit to the observed galaxy fluxes through a minimisation; this routine is extensively described in M. Sullivan et al. (2010) and, in the context of DES galaxy catalogues, M. Smith et al. (2020). From the best fitting templates, we adopt the associated parameters such as: redshift, stellar mass, and star formation rates throughout the rest of this work. The catalogue is complete to mag in -band, which is well-matched to the transient hosts that LSST should detect (e.g., P. Wiseman et al., 2020).
Host galaxies are associated to simulated SNe Ia following the observed rates as a function of the host-galaxy properties as measured by M. Sullivan et al. (2006). Within each galaxy, we follow the standard snana prescription of placing transients in the host following a Sérsic profile (J. L. Sérsic, 1963). We also ensure that the well-known correlation between the SN Ia parameter and host galaxy stellar mass is reproduced in the simulations (see section 4.5.1 in V21 for details), but we do not include the more subtle relation between SN Ia Hubble residuals and host galaxy properties, such as stellar mass, colour, and specific star formation rate.
For peculiar SNe Ia, we scale the SN Ia rate and assign host galaxies so that SN1991bg-like events are only associated to old stellar environments (e.g. W. Li et al., 2011; F. H. Panther et al., 2019) and SN Iax events are placed in star forming galaxies (e.g. J. D. Lyman et al., 2013, 2018; T. Takaro et al., 2020). For core-collapse SNe, we use the rates and host galaxy association measurements presented by O. Graur et al. (2017a, b). For the remaining transients their relation to host galaxies is not as extensively studied due to their rarer intrinsic and observational rates; we assign them to hosts randomly and position them within the host following a Sérsic profile. This random host association for the sub-dominant populations has no significant effect on the TiDES host galaxy statistics.
3.3 Simulating AGN
AGN simulations have been carried out based on properties of the selected sample. Due to the nature of optical selection in the target catalogues, the overwhelming majority of the sources will be broad line AGN, as desired for a reverberation mapping experiment. We identify each selected source with a spectroscopic template of a type 1 AGN as provided by the 4MOST AGN survey (A. Merloni et al., 2019). The associated template spectrum for each object has been redshifted and flux scaled according to the catalogue -band magnitude and fed into the 4MOST operations simulator. Within a 4MOST simulation, the AGN templates are used to determine the required exposure times in the DDFs to meet the spectral success criteria (Section 4.2).
3.4 Simulating 4MOST
Having generated a simulation of transient light curves as observed by LSST, we next determine which are available to be observed by 4MOST. The 4MOST observing strategy fulfils a complex balance of science requirements from multiple different surveys focused on both galactic and extra-galactic fields. To accommodate this, 4MOST will tile the entire southern-sky over the course of its five-year lifetime (R. S. de Jong et al., 2019), with the density of targets and exposure time requirements ultimately dictating the frequency of any returning visits. The overlap between the LSST wide-fast-deep footprint and the 4MOST union of observing tiles is shown in Fig. 2.
There are two important considerations in the apparent overlap between LSST and 4MOST. The first is that only a small fraction of live LSST transients will ever be observed by 4MOST due to the significant difference in the nightly sky coverage between the two surveys. The second is that although TiDES does have priority fibre-allocation on transient targets and their host galaxies in any field that 4MOST observes, it does not (and cannot) operate in a ‘Target of Opportunity’ mode and force observations of the most interesting targets by changing the telescope pointing. The exact pointing of 4MOST is not known more than a few days in advance, and can change according to predicted or actual observing conditions. Thus, at any one time, all possible LSST transient targets need to be available for insertion into the 4MOST queue, depending on the exact observing pointings that are chosen on a given night.
4MOST has experimented with several mock strategies to establish its observing plan. The success of any strategy is evaluated using a figure-of-merit for each 4MOST survey, combined into a global figure of merit. We evaluate the success of TiDES using the total number of objects observed and the quality of their resulting spectra. Our simulations in this paper use a typical 4MOST mock strategy that has been optimised to a solution that tiles the sky with observations (E. Tempel et al., 2020b) using probabilistic fibre assignments (E. Tempel et al., 2020a). From the TiDES perspective, we can use these simulations to track which of our simulated transients were observed and the exposure time on that epoch. For the host galaxies, we are able to assess the total integrated time the target received over multiple epochs.
3.5 LSST-TiDES selection function
We next describe the process of selecting transients from the LSST simulations for 4MOST spectroscopic follow-up. The aim is to reflect the ‘real-life’ operation of TiDES, where live SNe are identified through difference imaging and distributed in alert packets in the LSST data stream, we will identify suitable targets through custom filtering in community full-stream alert brokers e.g. Lasair (K. W. Smith et al., 2019), ALeRCE (F. Förster et al., 2021), ANTARES (T. Matheson et al., 2021), AMPEL (J. Nordin et al., 2019), FINK (A. Möller et al., 2021), Pitt-Google (W. M. Wood-Vasey, 2024), and Babamul. This means that new targets must be frequently added to the TiDES target list, and targets that have faded must be removed.
The depth of LSST is fainter than the spectroscopic limiting magnitude of 4MOST, allowing us to track the rising brightness of a transient before submitting it to the 4MOST observing queue – resulting in a high purity of candidates. Significant advances have been made in rapid real-time classification of transients and spectroscopic follow-up recommendation algorithms (e.g., D. Muthukrishna et al., 2019a; N. Kennamer et al., 2020; H. Qu & M. Sako, 2022; A. Gagliano et al., 2023; A. Möller et al., 2024; X. Sheng et al., 2024). However, TiDES has access to sufficient fibres to observe all live transients444In well-sampled LSST fields, live-transients in a 4MOST FoV pass our selection criteria in any field down to its limiting magnitude, and so is able to adopt a very straight forward selection of potential targets. This also has the advantage of a simpler calculation of the efficiency and selection function for any subsequent TiDES population analysis, including for SN Ia cosmology.
Our planned selection criteria for LSST transients, and those that we use in these simulations, are:
1.
The transient must be detected to in at least three bands over at least two nights,
2.
The object must be mag in any filter (Section 4.3 for justification).
We only consider data points, regardless of additional observations in of the field. Clearly, different filter selection criteria would be adopted for transients that did not originate from the LSST alert stream.
Once all criteria are met, the target is added to a mock TiDES observing queue with a nominal lifetime of four days, after which the target is retired. Each object, however, is monitored so that each subsequent LSST observation of the target that is brighter than 22.5 mag refreshes the target in the queue by extending its expiry date by an additional four days. This process ensures a high probability that any target selected by the 4MOST scheduler will be bright enough to have a successful spectroscopic observation (see our Spectral Success Criteria (SSC) in Section 4.2).
The requirements for the TiDES-Hosts programme are similar to that of TiDES-Live, but with no magnitude limit on the SN. We replace this with a requirement that the (now faded) SN light curve must have shown an increase in brightness followed by a decrease in brightness in at least two filters, followed by at least 50 days undetected in LSST data. This acts as a loose proxy for estimating when the transient peaked in brightness and faded below a detectable limit. We use this requirement as it ensures that the SNe in the selected hosts had both the pre- and post- maximum light data required for a reliable light-curve fit and photometric classification. This is needed for one of our primary science goals, constructing a SN Ia cosmology sample using the host-galaxy redshift. Additionally, we require that the host galaxy has , as targets fainter than this will never meet our SSC. Finally, TiDES-Hosts targets have no expiry date in the 4MOST queue.
It is possible to imagine different selection criteria and have several streams operating in parallel, each optimised towards transients with differing properties. Adapting some of our requirements may, for example, enable earlier triggering of young or rapid transients at the risk of increasing contamination from spurious events. The selection criteria presented in this work are the result of a conscious effort to maximise the quality of the targets we observe given our available fibre hours, but these may be revised once LSST and 4MOST are operational.
In the next section, we present the simulations, generate realistic spectra based on the observational metadata, and evaluate the expected samples.
4 Expected results from 4MOST/TiDES
The results from our simulations provide us with a catalogue of SNe, their hosts, and AGN that were ‘observed’ as part of the survey realisation. This includes metadata for the observational conditions, including the seeing, exposure time, and airmass. Combining all of this, we are able to generate mock spectra for each of our observed events, evaluate the quality of the mock spectra, and then quantify the performance of the survey simulation. We note that the conclusions we draw here are demonstrated at the population-level, and we do not consider any individual object assessment or seek to derive astrophysical properties from the spectra. Based on the density of targets and signal-to-noise requirements for each of our three science programmes, our fibre-hours are proportionally spent between TiDES-SN, TiDES-Hosts, and TiDES-RM respectively.
4.1 Mock spectra
To generate mock spectra, template spectra for each observed event are used as inputs to the 4MOST exposure-time calculator (ETC)555https://escience.aip.de/readthedocs/OpSys/etc/master/index.html, with parameters matching the observing conditions from the simulations. For the TiDES-Live survey, we generate phase-specific spectra based on the date of the 4MOST observation using the same templates described in 3.1. Where the same SN is observed more than once, we do not stack the spectra to improve the signal-to-noise due to evolution in the target spectrum. Conversely, for the TiDES-Hosts our targets can be repeatedly observed on multiple occasions and are not expected to evolve, and we take the total integrated exposure time over the survey when generating our final spectra. The ETC returns the instrument-level signal, noise, sky background, and the efficiencies for each of the three arms on the LRS.
An example spectrum for a peak-phase SN Ia, observed at different -band magnitudes, is shown in Fig. 3. The spectra for TiDES-Hosts (and TiDES-RM) are generated in an similar way.
4.2 Spectral success criteria
Once our targets have been simulated and observed by 4MOST, we use ‘Spectral Success Criteria’ (SSC) to determine whether an observation was successful. Each of the three surveys in TiDES has its own SSC. By design, objects that reach their SSC are not allocated any further fibre-hours on repeat visits to that field. This constraint is only relevant to the TiDES-Hosts survey as our other programs require the SSC to be achieved in a single visit.
For TiDES-Live, our actual success will depend on the signal-to-noise ratio (SNR) in the 4MOST spectrum and the transient type, due to the prominence of certain spectral features that define the different classes. SN spectra are typically dominated by broad features in the photospheric phases that are many tens-of-Angstroms wide, and thus our SSC are defined using the mean SNR in 15Å bins – we denote this with the SNR parameter. Our criteria are based on earlier studies of high-redshift SNe Ia (C. Balland et al., 2009), where a robust classification was achieved with a mean SNR = 5 over 4500–8000Å in the observer frame. This is a conservative criterion as C. Balland et al. (2009) also demonstrate probable classifications of transients with a mean SNR = 3.
We have checked our likely classification limits for SNe Ia using the ETC based on simulated mock spectra with classifications based upon the machine-learning SN classification tool dash (D. Muthukrishna et al., 2019b). dash can classify, with 95 per cent confidence, a SN Ia spectrum with SNR = 5 without using the host redshift as a prior. These simulations were performed with SN spectra free of host galaxy contamination (e.g., the SN Ia spectra in Fig. 3). An investigation into the effect of host galaxy contamination on SN classification in 4MOST will be presented in Milligan et al.(in prep).
For TiDES-Hosts, we require a mean SNR3 per Å over 4500–8000Å (SNR ) in the observer-frame. Galaxies with strong emission lines in real observations can have a redshift successfully measured before reaching this SNR criterion. In our simulation, however, we only obtain the final spectrum SNR rather than any intermediate stages that would have contributed to the stacked spectrum, we assume every galaxy reaching our SSC has had a redshift sucessfully measured.
For TiDES-RM, spectral success is defined by achieving a SNR =10 for an AGN spectrum. Just as for the TiDES-Live program, the spectra vary with time and, therefore, cannot be stacked if observations are taken on significantly different epochs. Within 4MOST, the other participating surveys have placed a high demand of time on the LSST DDFs. This necessitates many repeat visits that TiDES will exploit to routinely observe our AGN for a reverberation mapping experiment. To meet the SSC, the TiDES-RM targets must reach SNR =10 in a single visit every d.
4.3 Live supernova populations
Combining the mock spectra for all our targets and the SSC, we are able to construct an expected TiDES-Live sample. In many cases, we split our population statistics into groups for objects satisfying the SNR and criteria. As expected, LSST’s WFD survey program provides the bulk of our transients, per cent, due to the much larger footprint compared to the DDFs.
We begin with the most general overview metric for any wide-field sky-survey – the redshift distribution. Fig. 4 shows the redshift distribution split into the different transient sub-samples from our simulations. Our results show that we can expect to observe around 30 000 SNe with (SNR ) in total. Due to their intrinsic brightness, SNe Ia are the most common transient that TiDES will observe, totalling around 18 000 objects with a peak in the redshift distribution at z. This results in the largest homogeneously collected SN Ia sample from any spectroscopic survey. As expected, the SN1991bg-like and SN Iax under-luminous populations of SNe Ia are found at lower redshifts and are intrinsically rarer than the normal counterparts.
SNe II are intrinsically the most common of all SN sub-types (e.g., W. Li et al., 2011; O. Graur et al., 2017b) and this is reflected in the first two redshift bins of the histogram in Fig. 4. However, their lower luminosity results in a lower peak redshift for the sample at z. Stripped-envelope SNe (SESNe) also belong to the massive-star progenitor, core-collapse SN family, but occur at per cent of the intrinsic rate of SNe II (C. Frohmaier et al., 2021). In our sample, SESNe follow a similar observed trend with a peak redshift of z.
SLSNe-I form a class of the brightest observed SNe with a luminosity function ranging from (C. R. Angus et al., 2019). As a result, they are observable to the highest redshifts of any objects in our simulated sample, and their intrinsically long (and time dilated) light curves increases the opportunities for TiDES to obtain a spectrum. We have truncated the SLSN redshift distribution in Fig. 4 for visualisation purposes, but note that we expect to obtain SLSNe-I up to z.
Given the intrinsic luminosity, rates, and light curve evolution timescales of both Calcium-strong transients (CaST) and TDEs, it is unsurprising that these objects form the smallest populations of our samples. Furthermore, as these are some of the least well understood transient classes, any recent updates to their observed properties will not be reflected in our transient models from PLAsTiCC. For example, a slightly brighter TDE luminosity function than the one we adopt is presented in Y. Yao et al. (2023). Regardless, the TiDES sample of these transients will be a significant increase over current sample sizes.
We analyse the efficiency of our survey as a function of the target object’s magnitude in Fig. 5. We define the efficiency as the fraction of objects meeting the SSC once a fibre has been placed on them. To estimate these efficiencies, we fit the following sigmoid of the functional form (K. Sharon et al., 2007; S. A. Rodney et al., 2014):
where is the efficiency, is the maximum efficiency, is the magnitude of the object in the LSST -band at the time of observation, is the magnitude at which 50 per cent of the sample meet our SSC, and captures the exponential roll-off (smaller numbers describe a steeper drop).
For the SNR parameter, we find and for SNR , . In both cases, the maximum efficiency is and . Live transient objects are added to, or removed from, the 4MOST follow-up queue based, in part, on their observed LSST magnitudes. Using the result as our follow-up limit ensures that we minimise observations resulting in spectra with a SNR too low for a classification.
4.4 Observational efficiencies of type Ia supernovae
One of the main objectives of TiDES is to perform a SN Ia cosmology experiment, so we next examine various survey performance metrics for SNe Ia. TiDES will obtain classification spectra for around 13 000 SNe Ia with SNR (and 18 000 SNe Ia SNR ). Randomly selected example light curves are shown in Fig. 6 for the bands split into different redshift ranges of objects that were successfully observed by 4MOST.
The phase at which spectra are obtained is an important metric, as the ability to classify an object depends both on the phase and SNR of the spectrum. Objects observed near peak-brightness will have the highest SNR and classifiers will have large template libraries of historical observations to draw upon. In Fig. 7 we show the phases (measured relative to rest-frame -band peak) when the SNe Ia are triggered for follow-up (Section 3.5) and of their spectral observations. Most SNe are identified by LSST before they reach peak brightness, and then typically observed by 4MOST a few days after peak brightness. A break down of the triggering and spectrum phases for all transient types are presented in Figure 14.
In the violin plot of Fig. 8, we show the SNR metric in different redshift bins. We see broad distributions of SNR in the lowest redshift bins that narrow with increasing redshift and fall to lower mean SNR values. This is explained by the light curve evolution of our transients: at low redshifts the window-of-opportunity to observe a transient is large and a spectrum can be obtained over many phases and SN brightnesses and hence SNR . At higher-, objects are only observed around the peak of their light curves with a narrower range in brightnesses, and hence the SNR distribution narrows with a smaller mean value.
Finally, we quantify the efficiency of obtaining a good-quality spectrum as a function of redshift. This metric analyses the success rate of the spectra meeting an SSC criteria and is visualised in Fig. 9 (top panel). Objects at the lowest redshifts are almost always observed with SNR , as these objects will be the brightest in the sample. Naturally, our SNR curve falls away more rapidly, as spectra of this quality become increasingly more difficult to obtain in a fixed exposure time.
The lower panel of Fig. 9 shows the fraction of good-quality spectra for all SNe Ia that we trigger spectroscopic follow-up requests for following the selection function in Section 3.5. As expected, the fraction of all LSST transients that TiDES will observe is small, highlighting the number of transients that Rubin will discover.
4.5 TiDES host galaxies
TiDES-Hosts is focused on spectroscopy of SN host galaxies. At the most fundamental level we aim to obtain secure redshifts for these hosts but, as exposure depth is built up for some targets, the increasing signal-to-noise will enable us to measure additional properties (e.g., star formation rate, gas phase metallicity (J. Dumayne et al., 2023)).
In TiDES, the galaxy spectra can be obtained on multiple epochs and then stacked to increase the signal-to-noise until the SSC in Section 4.2 is reached. During 4MOST operations, automatic and manual redshifting will be performed on the new or stacked data products. Galaxies with prominent emission lines in their spectra can have their redshifts accurately determined before the continuum-defined SNR SSC are satisfied. A feedback-loop in the 4MOST system will remove such targets before additional fibre-hours are unnecessarily spent. However, this feedback is not incorporated in our simulations and instead observations are continued until the SSC is satisfied.
We generate galaxy mock spectra as described in Section 4.1 and apply the technique to every galaxy (Section 3.2) observed in the 4MOST-TiDES simulation. In total 400 000 galaxies were targeted by 4MOST but not all satisfied our requirement of SNR (Section 4.2) in the continuum. This is because the galaxies are injected during survey operations, resulting in fewer opportunities to add observations into a stacked spectrum as time progresses. In Figure 10, we show the redshift distribution split on both the total targeted by 4MOST those achieving SNR . Of these, our simulations show that 135 000 hosted SNe Ia. We assume in later sections of our analysis, that all objects reaching our SSC will have a reliable redshift measured from the host galaxy spectrum. Next, we demonstrate a cosmology analysis from our combined results of live SNe Ia (Section 4.3) and from their host galaxy spectra.
5 Supernova Cosmology
In this Section, we will present our expectations for the constraints TiDES can place on cosmological parameters following the methods employed by current generation experiments such as the Dark Energy Survey (DES Collaboration et al., 2024). For this analysis, the spectroscopic classification and redshift measurement is assumed to be perfect from the TiDES-Live program, and objects without live-spectra but with host spectroscopic redshifts are also assumed to have a perfect photometric typing based on the light curves.
The two SN Ia samples are based on the results of our LSST+TiDES simulations and are determined as follows:
‘LSST+TiDES photometric SN sample’: the sample of photometrically identified SNe Ia from their LSST light curves, for which the spectroscopic redshift of the SN host galaxy has been measured by TiDES following the SNR in the continuum cut.
Following the forecasts presented in the Dark Energy Science Collaboration (DESC) Science Requirement Document (The LSST Dark Energy Science Collaboration et al., 2018), we also include an external low- sample in our cosmology analysis of 2 400 SNe Ia at . The low- SNe in our analysis follows the simulations of the Foundation SN sample (R. J. Foley et al., 2018) presented by D. O. Jones et al. (2019). The expectation, however, is that this final sample size will only be achieved towards the end of 10-years of LSST operations, even though TiDES is expected to conclude after the first five years of LSST.
5.1 Defining the SN sample
We fit all simulated SNe Ia with the SALT2 light curve fitter (J. Guy et al., 2007) based on the simulated observed photometry. To ensure meaningful light curve fits, we follow standard requirements (e.g. D. Brout et al., 2019) of: i) observations in at least two different filters with SNR5, ii) at least one data point before the time of peak brightness, , and iii) at least one data point ten days after . The first of these are already satisfied by the the spectroscopic triggering criteria in Section 3.5, and the final two are satisfied in the majority of cases due to the high cadence of WFD in LSST.
Once the model has been fit to light curves, a final subset of the SNe Ia is selected by applying the following SALT2-based cuts (e.g., M. Betoule et al., 2014; D. Brout et al., 2019): , , , days, and fit probability 0.001. Those that pass this step are considered ‘cosmologically useful’ supernovae. These selection criteria are generally adopted in SN Ia cosmology analyses to reduce contamination from highly-reddened or other peculiar thermonuclear SNe while ensuring a minimal loss of ‘cosmologically useful’ SN Ia events. We find 12 600 successful SALT2 fits in the spectroscopic sample and 131 000 successful fits in the photometric SN sample. The redshift distribution for both samples and the low- sample are shown in the top panel of Fig. 11.
5.2 SN distance estimation and bias corrections
After light-curve fitting, we measure SN Ia distances and build the redshift–distance relation (‘Hubble diagram’) from which cosmological parameters can be constrained. We estimate the SN distance modulus , which is proportional to the logarithm of the SN luminosity distance, by applying the Tripp formula (R. Tripp, 1998)
where symbols are as defined in Section 3.1.1. The nuisance parameters , and are determined following the approach presented in J. Marriner et al. (2011). We present our final Hubble diagram in Fig. 11.
The term in Equation 3 is a bias correction term that is applied to each SN to correct for survey selection effects. In this work, bias corrections are measured using the BEAMS with Bias Corrections (BBC; R. Kessler et al., 2019b) code. BBC uses large simulations of the survey to model and correct for selection effects while simultaneously fitting for the nuisance parameters and intrinsic SN scatter. For our analysis, corrections for selection effects are modelled as a function of redshift only (so-called ‘1D’ corrections) and they are estimated using the same LSST+TiDES simulated samples used to measure the cosmological constraints.
The output of BBC is a redshift-binned Hubble diagram corrected for selection effects, and the associated diagonal covariance matrix, , that includes statistical uncertainties only for each redshift bin.
5.3 Cosmological Parameters
We use the output of the BBC fit and statistical+systematic covariance matrix to estimate cosmological contours for a FlatCDM model (i.e., assuming a constant value of ), and for a FlatCDM model, where and is the scale factor. We estimate cosmological contours using wfit.exe minimization code implemented in snana.
The cosmological contours from our simulations are presented in Fig. 12 and Fig. 13. We present constraints from the LSST+TiDES SN events (both spectroscopically-confirmed SNe Ia and SNe Ia with host spec- only) and the external low- sample. We compare our results with constraints from a simulation of the Dark Energy Survey 5 year SN sample (DES-SN5YR, The DES Collaboration, 2024). The DES-SN5YR analysis constitutes the state-of-the-art of SN cosmological measurements and provides some of the tightest constraints on the Dark Energy Equation of State. The DES-SN5YR simulation used in this work is generating assuming the same cosmology used in our LSST simulations (i.e., FlatCDM cosmology with ) and reflects the statistical power and data quality of the real DES-SN5YR data. While the published DES-SN5YR include statistical and systematic uncertainties, in Fig. 12 and Fig. 13 we only consider statistical uncertainties. For a FlatCDM model, LSST+TiDES SN samples combined with external low- samples and a CMB prior (Planck Collaboration et al., 2020) provide constrains on the Dark Energy Equation of State parameter, , of 0.012, 10 times smaller compared to DES-SN5YR. For a FlatCDM model, the Figure of Merit obtained from LSST+TiDES SN samples combined with external low- samples is 85, 15 times larger than the FoM of DES-SN5YR. When fitting for a FlatCDM model, we also combine SN data with a prior based on the recent Baryonic Acoustic Oscillation (BAO) measurements published by the Dark Energy Spectroscopic Instrument (DESI) Collaboration (A. G. Adame et al., 2025) using their first year of data. The DESI-BAO-Y1 prior is applied assuming the same true cosmology used to generate the LSST and DES-SN5YR simulations. With the DESI-BAO-Y1 prior, the LSST+TiDES FoM increases to 99 (7 times larger compared to DES-SN5YR combined with the same DESI-BAO-Y1 prior).
6 Summary
We have presented the Time-Domain Extragalactic Survey (TiDES) – a spectroscopic survey focused on understanding the extragalactic transient universe, and conducted on the 4m Multi-Object Spectroscopic Telescope (4MOST). TiDES will use 250 000 fibre-hours to address three key science goals: i) Create a Hubble Diagram of 140 000 cosmologically-useful SNe Ia to constrain cosmological measurements, including the dark-energy equation-of-state, , to sub-2 per cent precision; ii) Map the diversity of the transient Universe by obtaining more than 30 000 spectra of supernovae from LSST; iii) Perform a reverberation mapping experiment on 700–1 000 AGN in the high-cadence Deep-Drilling Fields.
In this work, we have simulated both the Rubin LSST and 4MOST surveys using the latest assumptions on their respective strategies. This has allowed us to create realistic light curves for millions of LSST transients with mock spectra for the sub-set that 4MOST can observe. Our results demonstrate that we will collate the largest homogeneous sample of supernovae and their host-galaxies to-date. Our results are summarised as follows:
We will obtain 30 000 SN spectra with a signal-to-noise ratio per 15Å (SNR ). Of these, 12 600 will be type Ia supernovae with a light curve quality suitable for a cosmological analysis.
TiDES will observe 9 000 core-collapse supernovae and create a sample of rare faint-and-fast transients larger than currently exists from any other survey to-date.
The TiDES-Hosts follow-up program will observe a sample of galaxies, of which will host SNe Ia suitable for cosmology.
Our cosmologically-useful sample, combining live-SNe Ia and photometrically identified SNe Ia with host galaxy redshifts, will be placed on a Hubble Diagram containing at-least 143 000 objects.
We performed a cosmology analysis of the LSST+TiDES sample, including a low- sample predicted by D. O. Jones et al. (2019). We are able to constrain the dark energy equations-of-state parameter, , to a sub-2 per cent statistical-only precision, this is 10 times smaller compared to the current gold-standard (DES Collaboration et al., 2024).
When we consider our constraints on a FlatCDM model, our LSST+TiDES SN sample produces a figure-of-merit value 15 times larger than DES-SN5YR.
The Rubin Observatory’s Legacy Survey of Space and Time will transform our understanding of the time-domain Universe. Important in unlocking the potential of these data will be the spectroscopic follow-up of transients, their host galaxies, and of active galactic nuclei. TiDES is primed to address this challenge providing dedicated follow-up time across the 5-year survey. All data collected by TiDES and other participating 4MOST surveys will be made public in periodic data releases on the ESO Science Archive Facility.
The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. {contribution} CF led the analysis, performed the simulations, and wrote the paper. MV and MSm performed the cosmology analysis. MSu discussed the results, implications, and commented on the manuscript at all stages of the project. SFH is PI of the TiDES-RM program and wrote the relevant sections of this paper. All other co-authors are members of 4MOST/TiDES and have commented on this work.
Appendix A Trigger and Spectrum Phases